OpenStudy (anonymous):

Suppose that 10 grams of Carbon is released in a nuclear energy plant accident. How long will it take for the 10 grams to decay to 1 gram? (Half-Life of Carbon (14C) is 5,715 years.

OpenStudy (mertsj):

@satellite73 http://openstudy.com/study#/updates/51356dfee4b093a1d949fdfb

OpenStudy (anonymous):

\[1=10\times 2^{\frac{t}{5715}}\] \[.1=2^{\frac{t}{5715}}\] then change of base formula \[\frac{t}{5715}=\frac{\ln(.1)}{\ln(2)}\]